OpenStudy (anonymous):

Find an equation of the line that (a) has the same y-intercept as the line y+2x-6=0 and (b) is parallel to the line -10x+11y=-4. Write your answer in the form y=mx+b.

OpenStudy (anonymous):

y + 2x -6 = 0 y = -2x + 6 (y-int when x=0 but since we rearranged it already we can see it is 6) y-int (0,6) parallel to the line -10x + 11y = -4 11y = 10x -4 y = 10/11x - 4/11 parallel meaning they have the same gradient we can now use the point (0,6) and gradient of 10/11 and sub it into the find an equation formula. y-y1 = m (x-x1) y - 6 = 10/11 (x-0) y - 6 = 10/11x y = 10/11 x + 6

OpenStudy (anonymous):

Thank you so much, this is all so confusing -_- lol

OpenStudy (anonymous):

no problem, it is actually quite easy after u understand the concept.